# RainPro-8: An Efficient Deep Learning Model to Estimate Rainfall Probabilities Over 8 Hours

### v1

@@ 3.4 Architecture
We design our architecture to forecast precipitation for a $512\times 512\textrm{km}^{2}$ patch. Assuming an average precipitation displacement rate of 1km per minute (Sønderby et al., 2020), we use 512km of spatial context on every side of the target patch ($\textrm{1536}^{2}\textrm{km}^{2}$) to provide our model with enough input context for all 8-hours of lead time.
Inspired by MetNet-3 (Andrychowicz et al., 2023), we use a U-Net (Ronneberger et al., 2015) with MaxViT (Tu et al., 2022) blocks to efficiently process multi-resolution data (Figure 1). Key differences include single-pass predictions without lead time conditioning (Section 3.3), early downsampling in the encoder, halving the internal channels, and removing topographical embeddings, all contributing to a reduced parameter count of 36.7M from the original 227M.
Figure 1: RainPro-8, optimized for reduced parameter count and improved efficiency, enabling simultaneous output of all timesteps and integration of data sources with varying resolutions.
The architecture begins with data fusion on the encoder, using Space-to-Depth convolutions (Sunkara and Luo, 2023) and ResNet blocks (He et al., 2016). Input data is transformed from $BTCHW$ to $B(TC)HW$ and merged with sources at matching resolutions. To balance efficiency and performance, we only use the full input context of 512km at 8km and 16km inputs, and use a lower input context of 256km at 4km resolution. The encoder aggregates features for each resolution into a low-resolution representation while preserving skip connections for the U-Net decoder. We apply downsampling before ResNet blocks instead of after to improve memory efficiency.
Next, the low-resolution representation is processed by 12 MaxViT blocks. MaxViT blocks, based on Vision Transformer (Dosovitskiy et al., 2021), combine local neighborhood and global gridded attention aggregate information across the full $\textrm{1536}^{2}\textrm{km}^{2}$ input patch. The decoder, using Transposed Convolutions (Long et al., 2015) and ResNet blocks, reconstructs high-resolution probability maps from the MaxViT output, while leveraging encoder skip connections. All forecast timesteps and intensity classes are generated by outputting $P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))$ (Section 3.2) in a single forward pass with timestamps encoded into the channel dimension (Section 3.3).
To ensure geographical alignment, representations are padded and cropped at each resolution stage. For example, 4km representations are padded equally on all sides before concatenating with 8km inputs, and the MaxViT output is cropped to the target region with extra context for upsampling, later cropped to $\textrm{512}^{2}\textrm{km}^{2}$ for the final output.
@@ 4 Experiments
RainPro-8 is implemented with PyTorch (Ansel et al., 2024) and PyTorch Lightning (Falcon and team, 2025), trained for 100k steps using a batch size of 16, validating every 2,000 steps. Training uses a static learning rate of 3e-4, AdamW optimizer with a weight decay of 0.1 and betas $(0.9,0.999)$, Exponential Moving Average (EMA) decay of 0.99975, dropout of 0.1, maximum stochastic depth (Huang et al., 2016) of 0.2, and lead-time decay rate of 10. We use 256 channels throughout the entire network, totaling 36.7 million parameters. Any other network hyperparameters follow those of MetNet-3 (Andrychowicz et al., 2023). The model with the lowest validation loss is selected. Training is performed on an NVIDIA H100 80GB SXM5 GPU and takes approx. 13 hours to complete.
@@
We use Critical Success Index (CSI) (Schaefer, 1990) at different thresholds, a standard accuracy metric in precipitation nowcasting (Andrychowicz et al., 2023; Gao et al., 2024b; Gong et al., 2024), and Continuous Ranked Probability Score (CRPS) (Hersbach, 2000), which assesses alignment of predicted probability distribution with observed values, rewarding sharp and reliable predictions. Fractions Skill Score (FSS) (Roberts and Lean, 2008) accounts for intensity shifts and offers tolerance to translation and deformation. Frequency Bias Index (FBI) (Termonia et al., 2018) quantifies over- and underforecasting, but not forecast quality. We also report Mean Absolute Error (MAE) and Mean Squared Error (MSE), but note that they are sensitive to high frequency of no-rain cases and thus less suitable for skewed distributions (Andrychowicz et al., 2023) (cf. App. B.1). Taking inspiration from the thresholding approach in MetNet (Espeholt et al., 2022), intensity-based metrics use the mean of the highest activated bucket of the cumulative output distribution. A bucket is activated if its predicted probability mass exceeds the corresponding threshold, where thresholds are computed for each bucket and lead time using the validation set. This approach effectively captures rare high-precipitation events, which tend to have lower predicted probability masses.
@@ 4.1 Performance Evaluation
We compare our model with global NWP GFS [^5] https://registry.opendata.aws/noaa-gfs-bdp-pds and extrapolation-based PySTEPS (Pulkkinen et al., 2019). PySTEPS extrapolates radar echoes for a few hours with a context of 512km. NWP forecasts are bilinearly interpolated to a 2 km/px resolution and temporally aligned to 10-minute intervals for radar comparison. Direct comparison with deep-learning nowcasting models is challenging due to fundamental limitations in their architectures and input assumptions. Specifically, models such as NowcastNet (Zhang et al., 2023), Earthformer (Gao et al., 2024b), and SimVP (Gao et al., 2022) operate only on radar data with fixed resolution and spatial coverage, and integrating additional data sources would require substantial architectural changes. To ensure a fair comparison, we provide all models with the same input region, including additional context, and evaluate them solely on the target central region. We train these models using their original architectures with MSE loss. SimVP, which forecasts sequences of the same length as its input, requires an autoregressive approach. Inputs are interpolated bilinearly from 4km to 2km for NowcastNet and SimVP. Earthformer, however, requires training at 4km resolution due to attention mechanism constraints, with outputs upsampled to 2km for evaluation. MetNet-3 (Andrychowicz et al., 2023) evaluation presents significant challenges due to its private code, reliance on US-specific data, and substantial computational requirements (227M parameters trained on 512 TPU v3 cores over 7 days). MetNet-3* is our faithful reimplementation of the architecture and training described in the original paper, adapted to our data and compute constraints, presenting a reproducible competitor under fair conditions. It incorporates lead-time conditioning and cross-entropy loss, and it only computes the loss on the high-resolution precipitation maps, not on accumulated rain or NWP initial state since we do not perform densification from weather stations.
As shown in Table 1, RainPro-8 outperforms all competitors on precipitation metrics — Critical Success Index (CSI) and Fractions Skill Score (FSS) — demonstrating strong skill and practical utility in rain forecasting. Frequency Bias Index (FBI) indicates that RainPro-8 slightly overpredicts precipitation, while PySTEPS is least biased. Although NowcastNet, SimVP, and Earthformer achieve lowest MAE and MSE due to their MSE loss in short-term nowcasting, these do not translate into higher CSI and FSS, likely because of domination of the no-rain class in the highly imbalanced dataset. RainPro-8 of performs slightly better than MetNet-3*, and offers additional advantages: a 48×48$\times$48 × inference speedup through single-step prediction and more coherent probability maps, enabled by the ordinal-consistent loss that accounts for class ordinality in probabilistic forecasts.
Table 1: Model comparison over all lead times and intensity thresholds.
@@
Figure 2 illustrates model performance (CSI) at different rain intensities and lead times up to eight hours (further plots and tables in App. B.2). Our model demonstrates superior performance across lead times and for all rainfall thresholds, ranging from light to heavy rain. The skill gap with extrapolation-based PySTEPS grows as lead times increase due to its assumption of constant motion and intensity. NowcastNet, Earthformer, and SimVP skill gap also grows with lead times due to the lack of additional data sources and probabilistic outputs. In contrast, GFS shows a larger skill gap at shorter lead times because of its longer convergence time. Performance is slightly lower for all models at higher precipitation intensities, likely due to the increased complexity in forecasting extreme weather events, which often involve more chaotic and unpredictable patterns.
@@
Figure 3 shows forecast for a $1000^{2}\textit{km}^{2}$ cropped region of at least 4 target patches of $512^{2}\textit{km}^{2}$, at three lead times (Europe-wide predictions visualized in App. D). Our model effectively captures most areas with rainfall, both in terms of intensity and location, with some blurriness as lead time increases due to growing uncertainty. GFS shows major deviations from the ground truth, with large differences in both intensity and rain-covered areas. PySTEPS, which can only extrapolate the latest radar image, struggles to capture changes in intensity or account for non-linear motion patterns at later lead times. NowcastNet, Earthformer, and SimVP are limited by inability to use more data sources—crucial for reliable 8-hour forecasts—and by MSE-loss, which favors frequent no-rain.
@@ 4.2 Ablation Studies
Table 2 shows ablation studies on the use of the ordinal-consistent loss function and single-pass predictions, compared to MetNet-3*, which uses non-ordinal cross-entropy loss and lead time conditioning (Andrychowicz et al., 2023). We also study the impact of two design choices in our loss function: class masking and lead time weighting, as well as the contribution of satellite data, topography, and GFS forecasts input. The results confirm the effectiveness of our approach, highlighting the advantages of the ordinal consistent loss with masking and single-pass predictions for generating 8-hour forecasts with lead time weights and all data sources. The slight gains of our model over the cross-entropy-based model are due to the latter enforcing ordinality only at inference time via cumulative probability computation, rather than learning it during training. Compared to models using lead-time conditioning, our single-pass prediction approach not only improves inference speed but also delivers superior performance across metrics. Excluding lead-time weights leads to a noticeable drop in performance, highlighting their importance in the loss. Similarly, the model without OC mask underperforms, likely because the model needs to focus on lower intensity classes before predicting higher ones, as identifying the first change in class activation (indicating expected rain) is more critical. Finally, utilizing all available data is key to achieving top forecasting accuracy.
Table 2: Ablation loss function, timesteps per forward pass, mask, lead time weights, data sources.
Model $\mathcal{L}$ Tsize OCmask OCltw Srcs. CRPS ()($\downarrow$)( ↓ ) CSI ()($\uparrow$)( ↑ ) FSS ()($\uparrow$)( ↑ )
RainPro-8 OC 48 $\surd$ $\surd$ $\surd$ 0.06096 0.2791 0.5367
MetNet-3* CE 1 - - $\surd$ 0.06199 0.2697 0.5173
RainPro-8_ce CE 48 - - $\surd$ 0.06098 0.2787 0.5357
RainPro-8_clt OC 1 $\surd$ - $\surd$ 0.06203 0.2695 0.5191
RainPro-8_noltw OC 48 $\surd$ $\times$ $\surd$ 0.06156 0.2733 0.5258
RainPro-8_nomask OC 48 $\times$ $\surd$ $\surd$ 0.06971 0.2565 0.5151
RainPro-8_sources OC 48 $\surd$ $\surd$ $\times$ 0.06331 0.2538 0.4910
@@ 4.3 Probability Maps
Our RainPro-8 model not only outperforms traditional operational methods in generating precipitation intensity maps, it also offers the advantage of quantifying uncertainty in predictions for each rainfall intensity. Figure 4 displays probability maps alongside ground truth for two different rain intensities and lead times, using the same sample and cropped region as in Figure 3, where intensity values were derived from these probability maps. The results show higher uncertainty for longer lead times and lower precipitation intensities, reflecting the inherent challenge of predicting scattered light rain far into the future at high resolution. For higher intensities, the model tends to predict low probabilities due to their low likelihood, particularly at extended lead times. Our model ensures consistency by guaranteeing that the probability of experiencing at least a certain precipitation level is always greater than or equal to the probability of experiencing higher precipitation levels.
@@ 4.4 Attribution
We study RainPro-8 using Integrated Gradients (IG) (Sundararajan et al., 2017), as done for MetNet-2 (Espeholt et al., 2022). IG attributes predictions to input features, offering insights into how the model utilizes each data source. Key IG findings (cf. App. C) are: high-resolution recent radar data is crucial for short lead times, while low-resolution radar gains importance around the 4-hour mark due to its broader coverage. Satellite data becomes important for lead times beyond 4 hours, where radar data is less effective. Visible channels have minimal impact, likely because they are restricted to daytime. GFS variables become more influential for longer lead times, with key contributors including wind components, vertical velocity, storm motion parameters, specific humidity, and surface metrics such as pressure and the Best Lifted Index (a measure of atmospheric instability), which are particularly relevant for forecasting cloud movement, precipitation, and thunderstorms. GFS forecasts become increasingly impactful at longer lead times, while less relevant for the first 4 hours.

### v2

@@ 3.4 Architecture
We design our architecture to forecast precipitation for a $512\times 512\textrm{km}^{2}$ patch. Assuming an average precipitation displacement rate of 1km per minute (Sønderby et al., 2020), we use 512km of spatial context on every side of the target patch ($\textrm{1536}^{2}\textrm{km}^{2}$) to provide our model with enough input context for all 8-hours of lead time. We use a U-Net (Ronneberger et al., 2015) with MaxViT (Tu et al., 2022) blocks to efficiently process multi-resolution data (Figure 1), similar to (Andrychowicz et al., 2023). Key differences include single-pass predictions without lead time conditioning (Section 3.3), early downsampling in the encoder, halving internal channels, and removing topographical embeddings, all contributing to a reduced parameter count of 36.7M from the original 227M.
Figure 1: RainPro-8 architecture, optimized for reduced parameter count and efficiency, integrating multiple data sources of varying resolution for simultaneous prediction of all lead times.
The architecture begins with data fusion on the encoder, using Space-to-Depth convolutions (Sunkara & Luo, 2023) and ResNet blocks (He et al., 2016). Input data is transformed from $BTCHW$ to $B(TC)HW$ and merged with sources at matching resolutions. To balance efficiency and performance, we only use the full input context of 512km at 8km and 16km inputs, and lower input context of 256km at 4km resolution. The encoder aggregates features for each resolution into a low-resolution representation while preserving skip connections for the U-Net decoder. To improve memory efficiency, we apply downsampling before ResNet blocks instead of after.
The low-resolution representation is processed by 12 MaxViT blocks. Based on Vision Transformer (Dosovitskiy et al., 2021), they combine local neighborhood and global gridded attention aggregate information across the full $\textrm{1536}^{2}\textrm{km}^{2}$ input patch. The decoder, using Transposed Convolutions (Long et al., 2015) and ResNet blocks, reconstructs high-resolution probability maps from the MaxViT output, while leveraging encoder skip connections. All forecast timesteps and intensity classes are generated by outputting $P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))$ (Section 3.2) in a single forward pass with timestamps encoded into the channel dimension (Section 3.3). For geographical alignment, representations are padded and cropped at each resolution. E.g., 4km representations are padded equally on all sides before concatenating with 8km inputs, and MaxViT output is cropped to the target region with extra context for upsampling, later cropped to $\textrm{512}^{2}\textrm{km}^{2}$ for final output.
@@ 4 Experiments
RainPro-8 is implemented with PyTorch (Ansel et al., 2024) and PyTorch Lightning (Falcon & team, 2025), trained for 100k steps using a batch size of 16, validating every 2,000 steps. Training uses a static learning rate of 3e-4, AdamW optimizer with a weight decay of 0.1 and betas $(0.9,0.999)$, Exponential Moving Average (EMA) decay of 0.99975, dropout of 0.1, maximum stochastic depth (Huang et al., 2016) of 0.2, and lead time decay rate of 10. We use 256 channels throughout the entire network, totaling 36.7 million parameters. Any other network hyperparameters follow those of MetNet-3 (Andrychowicz et al., 2023). The model with the lowest validation loss is selected. Training is performed on an NVIDIA H100 SXM5 GPU and takes approx. 13 hours (cf. App. F).
@@
We use Critical Success Index (CSI) (Schaefer, 1990) at different thresholds, a standard accuracy metric in precipitation nowcasting (Andrychowicz et al., 2023; Gao et al., 2024b; Gong et al., 2024), and Continuous Ranked Probability Score (CRPS) (Hersbach, 2000), which assesses alignment of predicted probability distribution with observed values, rewarding sharp and reliable predictions. Fractions Skill Score (FSS) (Roberts & Lean, 2008) accounts for intensity shifts and offers tolerance to translation and deformation. Frequency Bias Index (FBI) (Termonia et al., 2018) quantifies over- and underforecasting, but not forecast quality. We also report Mean Absolute Error (MAE) and Mean Squared Error (MSE), but note that they are sensitive to high frequency of no-rain cases and thus less suitable for skewed distributions (Andrychowicz et al., 2023) (cf. App. B.1). Taking inspiration from the thresholding approach in MetNet (Espeholt et al., 2022), intensity-based metrics use the mean of the highest activated bucket of the cumulative output distribution. A bucket is activated if its predicted probability mass exceeds the corresponding threshold, where thresholds are computed for each bucket and lead time using the validation set. This approach effectively captures rare high-precipitation events, which tend to have lower predicted probability masses.
@@ 4.1 Performance Evaluation
RainPro-8 is benchmarked against global NWP GFS [^5] https://registry.opendata.aws/noaa-gfs-bdp-pds, and PySTEPS (Pulkkinen et al., 2019) which extrapolates radar echoes with a 512 km context. GFS forecasts are bilinearly interpolated to 2 km/px and temporally aligned to 10 min for radar comparison. No other deep-learning nowcasting model can learn from all our input sources: Earthformer (Gao et al., 2024b) and SimVP (Gao et al., 2022) operate only on radar with fixed resolution and coverage, and other data sources cannot be handled by their architectures. For fairness, all models use the same input region (with added context) but are evaluated only on the central target region, including a radar-only RainPro-8 variant (RainPro-8R) to demonstrate that our approach provides further performance benefits beyond its multiple data source capabilities. SimVP, which produces outputs matching its input length, is run autoregressively with inputs interpolated from 4 km to 2 km to achieve the target resolution. Earthformer is trained at 4 km due to attention limits, with outputs upsampled to 2 km for evaluation. MetNet-3 (Andrychowicz et al., 2023) is challenging to evaluate due to its private code, reliance on US-specific data, and substantial computational requirements (227M parameters trained on 512 TPU v3 cores over 7 days). MetNet-3* is our faithful reimplementation of the architecture and training described in the original paper, adapted to our data and compute constraints, presenting a reproducible competitor under fair conditions. It incorporates lead time conditioning and cross-entropy loss, and it only computes the loss on the high-resolution precipitation maps, not on accumulated rain or NWP initial state since we do not perform densification from weather stations.
As shown in Table 1, RainPro-8 outperforms all competitors on both precipitation metrics, CSI and FSS. RainPro-8 performs slightly better than MetNet-3*, and offers additional advantages: a $48\times$ inference speedup through single-step prediction and more coherent probability maps, enabled by the ordinal-consistent loss that accounts for class ordinality in probabilistic forecasts. RainPro-8R surpasses existing radar-only baselines, demonstrating that our approach offers superior performance, but also lagging behind RainPro-8 at longer lead times, underscoring the benefit of our multi-source integration in our full model. FBI indicates that RainPro-8 slightly overpredicts precipitation, Earthformer and SimVP significantly underpredict, and PySTEPS is the least biased. Although SimVP and Earthformer achieve lowest MAE and MSE due to their MSE loss, these do not translate into higher CSI or FSS, likely suffering from impact of the dominating no-rain class.
Table 1: Performance (CSI, FSS), bias (FBI), error (MAE, MSE) across lead times and intensities.
@@
Figure 2 illustrates model performance (CSI) at different rain intensities and lead times up to eight hours (further details in App. B.2). Our model demonstrates superior performance across lead times and rainfall thresholds, from light to heavy rain. The skill gap with extrapolation-based PySTEPS grows as lead times increase due to its assumption of constant motion and intensity. The skill gap also widens with lead time, where Earthformer and SimVP lack sufficient input and probabilistic output capabilities. RainPro-8R confirms the impact of lack of additional data sources over time. In contrast, GFS shows a skill gap at shorter lead times because of its longer convergence time. Performance is slightly lower for all models at higher precipitation intensities, likely due to complexity in forecasting extreme weather, which often involves more chaotic and unpredictable patterns.
@@
Figure 3 visualizes a $1000^{2}\textit{km}^{2}$ cropped region of at least 4 target patches of $512^{2}\textit{km}^{2}$, at 3 lead times (Europe-wide in App. D). RainPro-8 effectively captures areas with rainfall, both in intensity and location, with some blurriness as lead time increases uncertainty. GFS shows major deviations from ground truth, with large differences in intensity and rain-covered areas. PySTEPS struggles to capture changes in intensity or account for non-linear motion patterns at later lead times as it only extrapolates the latest radar image. RainPro-8R, Earthformer, and SimVP cannot exploit additional data, and the latter two suffer from MSE loss favoring no-rain events.
@@ 4.2 Ablation Studies
Table 2 shows ablation studies of our ordinal-consistent loss, single-pass predictions, lead time weighting and multi-source input, compared to MetNet-3* using non-ordinal cross-entropy loss and lead time conditioning (Andrychowicz et al., 2023). Results confirm the effectiveness of our approach, and the contribution of each of our design choices and learning strategies. The slight yet consistent gains of our model over the cross-entropy-based model (cf. App G) are due to the latter enforcing ordinality only at inference time via cumulative probability computation, rather than learning it during training. Compared to models using lead time conditioning, our single-pass prediction not only improves inference speed (cf. App F) but also delivers superior performance across metrics. Excluding lead time weights leads to a noticeable drop in performance. Finally, an architecture capable of integrating all available data is key to top accuracy compared to radar-only RainPro-8R.
Table 2: Ablation loss function, timesteps per forward pass, lead time weights, data sources.
Model $\mathcal{L}$ Tsize OCltw Srcs. CRPS $(\downarrow)$ CSI $(\uparrow)$ FSS $(\uparrow)$
RainPro-8 (ours) OC 48 $\surd$ $\surd$ 0.06096 0.2791 0.5367
CE loss (no OC) CE 48 $\surd$ $\surd$ 0.06098 0.2787 0.5357
Lead time conditioning OC 1 - $\surd$ 0.06203 0.2695 0.5191
No lead time weights OC 48 $\times$ $\surd$ 0.06156 0.2733 0.5258
Radar-only (RainPro-8R) OC 48 $\surd$ $\times$ 0.06574 0.2289 0.4491
MetNet-3* CE 1 - $\surd$ 0.06199 0.2697 0.5173
@@ 4.3 Probability Maps
RainPro-8 not only outperforms baselines in precipitation intensity forecasts but also quantifies uncertainty for each intensity. Figure 4 displays probability maps and ground truth for two rain intensities and lead times (same sample and crop as Figure 3), with intensities derived from these probabilities. Uncertainty is higher for longer lead times and light rain, reflecting the challenge of predicting scattered light rain far into the future at high resolution. Higher intensities are predicted with lower probabilities because of their low likelihood, particularly at extended lead times. The model ensures consistency by never assigning higher probabilities to higher precipitation levels.
@@ 4.4 Attribution
We analyze RainPro-8 with Integrated Gradients (IG) (Sundararajan et al., 2017), as in MetNet-2 (Espeholt et al., 2022), attributing predictions to inputs to reveal how data sources are used. Rather than isolating each source with separate experiments, we use IG to assess their relative importance across lead times and variables (cf. App. C). Results show recent high-res radar drives short lead times, while low-res radar gains value around 4 hours for broader coverage. Beyond 4 hours, satellite data becomes more useful, though visible channels contribute little due to daytime limits. GFS variables become more influential for longer lead times, with key contributors including wind components, vertical velocity, storm motion parameters, specific humidity, and surface metrics such as pressure and the Best Lifted Index (a measure of atmospheric instability), which are particularly relevant for forecasting cloud movement, precipitation, and thunderstorms. GFS forecasts become increasingly impactful at longer lead times, while less relevant for the first 4 hours.
4.5 Radar-Only Short-Term Nowcasting (SEVIR Benchmark)
While RainPro-8 targets 8-hour high-resolution precipitation forecasts using multi-source data, we also evaluate a radar-only short-term version (RainPro-2R) for 2-hour predictions on the SEVIR benchmark for direct comparison with existing nowcasting models (App. H). Table 3 shows RainPro-2R outperforms state-of-the-art deterministic models. Compared to generative models, it achieves higher CSI and HSS (pixel-wise metrics) but lower pooled CSI and perceptual metrics (LPIPS, SSIM), reported in Yu et al. (2024) (App. H). Pooled CSI only checks whether any pixel within a neighborhood is hit, ignoring how many pixels match, whereas FSS rewards both correct pixels and the fraction of pixels correctly predicted, giving a more nuanced measure of spatial accuracy at full resolution (Figure 5). Perceptual metrics reflect a trade-off: sharper outputs sacrifice per-pixel performance (Yan et al., 2024). RainPro-2R’s probabilistic predictions appear blurrier but capture uncertainty and higher accuracy. RainPro-2R also outperforms DiffCast on CRPS and FSS while being 13× faster at inference. These results highlight RainPro’s performance both for the targeted 8-hour multi-source forecasting and SEVIR’s short-term radar-only nowcasting setting.
Table 3: SEVIR benchmark; best results in bold per category; from DiffCast (Yu et al., 2024)
Method CSI $(\uparrow)$ CSI-p4 $(\uparrow)$ CSI-p16 $(\uparrow)$ HSS $(\uparrow)$ LPIPS $(\downarrow)$ SSIM()($\uparrow$)
Deterministic
RainPro-2R 0.3524 0.3834 0.4171 0.4501 0.2353 0.5966
PhyDnet 0.2560 0.2685 0.3005 0.3124 0.3785 0.6764
MAU 0.2463 0.2566 0.2861 0.3004 0.3933 0.6361
ConvGRU 0.2416 0.2554 0.3050 0.2834 0.3766 0.6532
SimVP 0.2662 0.2844 0.3452 0.3369 0.3914 0.6304
Earthformer 0.2513 0.2617 0.2910 0.3073 0.4140 0.6773
Generative
STRPM 0.2512 0.3243 0.4959 0.3277 0.2577 0.6513
MCVD 0.2148 0.3020 0.4706 0.2743 0.2170 0.5265
PreDiff 0.2304 0.3041 0.4028 0.2986 0.2851 0.5185
DiffCast 0.3077 0.4122 0.5683 0.4033 0.1812 0.6354
Figure 5: CRPS and FSS across different neighborhoods, thresholds, and lead times.